OpenStudy (anonymous):

find the average rate of change of f(x)=tan x from 0 to pi/4

OpenStudy (anonymous):

rate of change: use this equation \[\frac{ f(b)-f(a) }{ b-a }\] where in your case f(x)=tanx and a=0 and b=pi/4